Question 4

A vehicle has a 20-gal tank and gets 15mpg. The number of miles N that can be driven depends on the amount of gas đť‘Ą in the tank:

a) Write a formula that models this situation.

b) Determine the number of miles the vehicle can travel on (i) a full tank of gas, and (ii)

3/4 of a tank of gas.

c) Determine the domain and range of the function.

d) Determine how many times the driver had to stop for gas if she has driven a total of

578 mile


1
Expert's answer
2021-09-21T05:02:45-0400

a) The model that describes this situation will be


"N{(x)}=(20\,gal)(15\frac{miles}{gal})x=[300\,miles]x"


b) we proceed to calculate the mileage by substituting x=1 and x=3/4 for the tank level:


"N(x=1)=[300\,miles](1)=300\,miles
\\ N(x=3/4)=[300\,miles](3/4)=225\,miles"


c) The domain of N(x) would be "0\leq x \leq 1" and the range will be "0\,miles\leq N(x) \leq 300\,miles"


d) We have to calculate x:


"N'=578\,miles=[300\,miles]x'
\\ \implies x'=578\,miles/300\,miles=1.926"


In conclusion, the driver has to stop only once during the whole trip to fill the tank.